OpenStudy (anonymous):

Use a calculator to evaluate cot π5. 0.561 1.376 91.185 32.142

OpenStudy (shamim):

Is it Cot5*pi=?

zepdrix (zepdrix):

It's asking you to use a calculator... what do you need help with? do you not have access to a calculator? need assistance inputting the expression?

OpenStudy (anonymous):

I have no access to a calculator! Its the cot of pi/5

OpenStudy (anonymous):

\[\cot \left(\frac{180 {}^{\circ}}{5}\right)=\sqrt{1+\frac{2}{\sqrt{5}}}=1.37638 \]